OpenStudy (anonymous):

PLEASEEEE HELPP!!!!!! If the arc on a particular circle has a length of 12 units, and the circumference of the circle is 36, which is the degree measure of the arc?

OpenStudy (anonymous):

tony hawk

jimthompson5910 (jim_thompson5910):

The arc is 12/36 = 1/3 of the full circumference that means the central angle is 1/3 of a full 360 degree turn so the central angle is (1/3)*360 = 120 degrees The arc measure is equal to the central angle. So the arc measure is also 120 degrees.

OpenStudy (anonymous):

thanks man

jimthompson5910 (jim_thompson5910):

yw